OpenStudy (anonymous):

Where can one find a proof or explanation of the equivalence posted via equation editor.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \left(1+\frac{a}{x} \right)^x=e^a\]

OpenStudy (anonymous):

That's really just the definition of e. http://mathworld.wolfram.com/e.html